OpenStudy (anonymous):

A country's population in 1994 was 140 million. In 1998 it was 146 million. Estimate the popluation in 2011 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (e.mccormick):

oops... wrong button... my bad.

OpenStudy (e.mccormick):

Well, since I am here... Do you know the exponential growth formula?

OpenStudy (here_to_help15):

. A = 140 P(4) = 146 = 140e^(4k) e^4k = 146/140 4k = ln (146/140) k = ln(146/140)/4 = 0.010491

OpenStudy (here_to_help15):

P(17) = 140e^(0.010491*17)=........

OpenStudy (anonymous):

Okay? Now?

OpenStudy (e.mccormick):

Put it in a calculator or use an approcimation for e.

OpenStudy (anonymous):

it keeps on saying error. :/

OpenStudy (e.mccormick):

Then you are putting something in wrong or it does not know what e is.

OpenStudy (anonymous):

p(17)=140e^.178347?

OpenStudy (anonymous):

Im, sorry this is a timed review, can someone please give me a direct asnwer?

OpenStudy (e.mccormick):

OpenStudy is for study, not timed evaluations.

OpenStudy (anonymous):

Yes, but this is the one question I need to pass, so please??

OpenStudy (anonymous):

:(

OpenStudy (here_to_help15):

I would but you heard the boss ^_^

OpenStudy (anonymous):

Lol, okay fine. Walk me through it, I'll give you a medal?

OpenStudy (here_to_help15):

@e.mccormick would you mind i have my own question to solve thanks :D

OpenStudy (anonymous):

Okay nevermind, you can do your own thing.